OpenStudy (anonymous):

A family decides to put tiles in the entryway of their home. The entryway has an area of 3 square meters. If each tile is 2 centimeters by 2 centimeters, how many tiles will it take to cover the entryway? It'll take_____ tiles to cover the entryway.

OpenStudy (anonymous):

@DaWizjr

OpenStudy (anonymous):

1 meter = 100 centimeters

OpenStudy (anonymous):

how would i set the problem up?

OpenStudy (anonymous):

hold on I will be back on later

OpenStudy (anonymous):

@undeadknight26 help her right quick

OpenStudy (anonymous):

@robtobey

undeadknight26 (undeadknight26):

Riiighhhhttt quiiicckkk XD I love those words for some reason lol

OpenStudy (anonymous):

lol! can u help me quick?

undeadknight26 (undeadknight26):

How many centimeters in a inch?

OpenStudy (anonymous):

2.54

undeadknight26 (undeadknight26):

Oh crude its meters and as @DaWizjr said: 1 meter = 1000 cm

OpenStudy (anonymous):

okay, so how do I set the problem up?

undeadknight26 (undeadknight26):

Well we know that 2x2 = 4

undeadknight26 (undeadknight26):

And we know that we need to cover 300cm so i would do 4/300 = .0133 Then i believe you would do 3/.0133 = 225

OpenStudy (anonymous):

hmm, i typed in answer and 225 is not correct

OpenStudy (anonymous):

unless we do 2.54*4?

undeadknight26 (undeadknight26):

hmm?

OpenStudy (anonymous):

@zimmah would u happen to know what we do here?

OpenStudy (anonymous):

the answer was 7500 @undeadknight26 so how do u think they got that answer?

undeadknight26 (undeadknight26):

@amistre64 Sorry to give you more work but i think i confused myself heres...

undeadknight26 (undeadknight26):

lol no idea xD

OpenStudy (anonymous):

okay, thnkyou

undeadknight26 (undeadknight26):

np

OpenStudy (anonymous):

@blurbendy

undeadknight26 (undeadknight26):

Sorry i haven't done this stuff in a while :/

OpenStudy (anonymous):

its alright:)

OpenStudy (blurbendy):

hey, you're back @Katlin95!

OpenStudy (anonymous):

yes!!

OpenStudy (anonymous):

can ya help me with a couple questions?

OpenStudy (anonymous):

I ws playing some games, but i can answer this

OpenStudy (anonymous):

 Let's say this large square has sides of 4m (so the area would be 4*4=16 m^2) And let's now divide it in 4 pieces.  Each piece has sides of 2m so an area of 2*2=4m^2 Of course the total area is 4+4+4+4 or simply 4*4)=16 m^2